A stream flows from A to B, a distance of 30 km in 2 km/hr. and a man can go up and down in 8 hours. what is the rate of man in still water.
Answer by malaydassharma(59)

Let the rate of man in still water = x Km/Hr
Hence, downstream speed of themn = (x+2) km/hr and upstream is (x-2) km/hr
As per the problem
30/(x+2) + 30/(x-2)=8
or 30(x-2)+ 30(x+2)=8(x+2)(x-2)
or 60x=8x^2-32
or 8x^2-60x-32=0
or 2x^2-15x-8=0
or (x-8)(2x-1)=0
hence, x=8 or 1/2 < 2 (speed of the river stream), which is not possible.
Hence, Rate of man in still water =8 km/hr